# DiTAR: Diffusion Transformer Autoregressive Modeling for Speech Generation

### v2

@@ Abstract
Several recent studies have attempted to autoregressively generate continuous speech representations without discrete speech tokens by combining diffusion and autoregressive models, yet they often face challenges with excessive computational loads or suboptimal outcomes. In this work, we propose Diffusion Transformer Autoregressive Modeling (DiTAR), a patch-based autoregressive framework combining a language model with a diffusion transformer. This approach significantly enhances the efficacy of autoregressive models for continuous tokens and reduces computational demands. DiTAR utilizes a divide-and-conquer strategy for patch generation, where the language model processes aggregated patch embeddings and the diffusion transformer subsequently generates the next patch based on the output of the language model. For inference, we propose defining temperature as the time point of introducing noise during the reverse diffusion ODE to balance diversity and determinism. We also show in the extensive scaling analysis that DiTAR has superb scalability. In zero-shot speech generation, DiTAR achieves state-of-the-art performance in robustness, speaker similarity, and naturalness.
@@ 1 Introduction
Recently, the autoregressive language model (LM) has demonstrated strong generative capabilities and scaling properties(Achiam et al., 2023; Touvron et al., 2023; Yang et al., 2024; Team et al., 2023), where discrete tokenization is commonly used. While discretization is natural for text, it is not as straightforward for other modalities, such as images, video, and audio. Due to bitrate limitations, discrete representation often fails to reconstruct complex modalities with high fidelity. In comparison, continuous tokens can better reconstruct the original information(Fan et al., 2024; Rombach et al., 2022). To alleviate this challenge, a two-stage strategy is commonly applied for autoregressive multi-media generation, where a lossy discrete token is first generated by an LM, followed by a token-based diffusion for detail enrichment. However, such a cascaded design suffers from error accumulation and limits the scalability of the language model. Therefore, autoregressive modeling of continuous representations is meaningful.
@@
The diffusion model has been proven effective at modeling continuous representations(Peebles & Xie, 2023; Rombach et al., 2022; Anastassiou et al., 2024), and integrating it with autoregressive models is a major research trend. (Li et al., 2024a) proposes incorporating a diffusion head into a language model for image generation, pioneering a new approach. However, the performance significantly falls short when a causal-attention language model forms the backbone. Another approach, such as ARDiT(Liu et al., 2024b) or Transfusion(Zhou et al., 2024), repurposes the language model’s parameters for diffusion, leading to substantial computational demands. This raises the question of whether we can autoregressively predict continuous tokens while ensuring high-quality generation and maintaining reasonable computational demands.
@@
To tackle this challenge, we introduce the Diffusion Transformer AutoRegressive (DiTAR) modeling, a framework that seamlessly combines transformer diffusion with a language model. We attribute the subpar performance of a language model with a diffusion head to the unidirectional dependency imposed by causal attention, which conflicts with the close inter-frame correlations characteristic of continuous tokens. In DiTAR, we propose a divide-and-conquer strategy that breaks continuous tokens into multiple patches. A language model is responsible for inter-patch prediction, while a diffusion transformer handles intra-patch prediction. Specifically, a diffusion transformer with bidirectional attention (DiT), noted for its state-of-the-art performance across various generative fields(Liu et al., 2024a; Peebles & Xie, 2023; Anastassiou et al., 2024), is employed to predict localized patches, named LocDiT. Furthermore, we propose a broadly applicable guidance method and introduce historical contexts to LocDiT to further enhance its generative capabilities. After patchification, the causal language model handling long contexts receives shorter sequences, further reducing computational load.
@@
For inference, temperature-based sampling balances exploration and exploitation and is crucial in discrete-valued language models. However, Its application in continuous-valued LMs remains underexplored. We define temperature as the point at which noise is introduced along the reverse ODE trajectory and propose a temperature-based sampling method for the continuous-valued AR model. Distinct from previous SDE-based approaches(Dhariwal & Nichol, 2021; Li et al., 2024a) that require numerous steps, our method suits quicker ODE solvers and adeptly balances diversity with determinism.
@@
DiTAR, rooted in a language model, excels at zero-shot generation tasks. We apply DiTAR to zero-shot text-to-speech (TTS) that aims to generate speech from unseen speakers’ prompts. Unlike systems(Chen et al., 2024b, 2025; Anastassiou et al., 2024; Du et al., 2024) using a coarse-to-fine pipeline, DiTAR simplifies the process by having the language model directly predict final features, obviating the need for multiple stages. It achieves state-of-the-art results in robustness, speaker similarity, and naturalness while demanding far less computational power than competing models.
@@ 2 Related Works
Integrating Autoregressive Language Model and Diffusion. Language models are primarily used for discrete representations, while diffusion excels in modeling continuous distributions. Integrating them for multimodal modeling is a crucial research direction. Some efforts (Wu et al., 2023; Liu et al., 2024b; Chen et al., 2024a) enable diffusion to have autoregressive capabilities by varying the denoising rates between consecutive tokens to achieve earlier predictions for preceding tokens. Transfusion(Zhou et al., 2024) utilizes a shared transformer for both diffusion and language models, employing causal attention for discrete tokens and bidirectional attention for continuous tokens. However, it does not naturally support the autoregressive generation of continuous tokens. These approaches repurpose language model parameters for diffusion, which significantly increases computational demands as the sequence lengthens and the language model size grows. Most relevant to our work, Li et al. (2024a) proposes a diffusion head for next-token prediction. However, its application in causal language models results in relatively poor performance.
@@
Patchification in Generative Modeling. In speech(Wang et al., 2017; Meng et al., 2024; Chen et al., 2024b), image(Peebles & Xie, 2023; Li et al., 2024b), and video generation(Liu et al., 2024a), the patchification technique is widely applied. In these works, patchification primarily aims to reduce the computational load by shortening the sequence length. However, in this paper, patchification not only lowers computational demands but also enables bidirectional modeling on patches within our autoregressive framework, further improving modeling effectiveness.
@@
Zero-Shot Text-to-Speech. Zero-shot TTS aims to generate speech with prompts from unseen speakers. Existing works can be divided into two categories: multi-stage and single-stage. Multi-stage approaches(Chen et al., 2024b, 2025; Lee et al., 2023; Jiang et al., 2023b, a; Anastassiou et al., 2024), represent speech using various representations, primarily divided into coarse and fine categories. The autoregressive language model is often adopted to predict the coarse representations, usually in discrete value and low information, such as semantics(Kharitonov et al., 2023; Anastassiou et al., 2024; Hsu et al., 2021; Chung et al., 2021; Baevski et al., 2020) and prosody(Jiang et al., 2023b, a; Ju et al., 2024). And then another model is adopted to conduct coarse-to-fine. Single-stage methods focus on generating high-information continuous representations, such as Mel spectrograms or latents of auto-encoders, which can directly reconstruct audio waveforms. Diffusion based on the non-causal attention transformer is the primary method employed(Chen et al., 2024c; Eskimez et al., 2024; Gao et al., 2023; Le et al., 2024). Additionally, some approaches(Meng et al., 2024; Wang et al., 2017; Shen et al., 2018; Li et al., 2019) directly use AR to model Mel-spectrograms of speech but often involve using Dropout(Srivastava et al., 2014) at the input stage to guarantee generation robustness, resulting in weaker in-context learning capabilities. These methods are better suited for in-set speaker TTS. This paper introduces a single-stage autoregressive method for zero-shot TTS achieving SOTA generation robustness and voice similarity.
3 Method
@@ 3.1.1 Formulation
DiTAR is an autoregressive model via next-token prediction. Consider a sequence of continuous tokens $\bm{x}=(\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{N})$, we can factorize the joint distribution of the sequence by the chain rule of probability:
@@
We validate the effectiveness of DiTAR on the zero-shot text-to-speech task. Based on the formulation, we regard zero-shot TTS as a conditional continuation task for the AR model like(Chen et al., 2024b), where prompting texts, target text, and prompting speech are concatenated and fed into the model as prefix context, then the model autoregressively generates the target speech given the context.
@@ 3.1.2 Overall architecture
Li et al. (2024a) finds that the performance of the causal-attention autoregressive model combined with diffusion loss is significantly inferior to systems utilizing full attention. To address the issue, we propose a divide-and-conquer strategy, where a long sequence of continuous tokens is divided into multiple patches. A language model is responsible for inter-patch prediction, while a diffusion transformer handles intra-patch prediction. As shown in Figure 1, the backbone of our system is a causal-attention transformer with next-token prediction. Each patch of continuous tokens is processed with an aggregation encoder into a single vector, which is then fed into the AR model to get the output embedding $\bm{h}_{t}$. $\bm{h}_{t}$ serves as the condition of the following diffusion decoder, LocDiT. Following (Li et al., 2024a), a diffusion loss is used for the output continuous tokens at training time.
@@ 3.2 LocDiT: Next-Patch Bidirectional Modeling
The diffusion transformer(Peebles & Xie, 2023; Liu et al., 2024a) has achieved success across numerous generative fields. These approaches leverage the full receptive field of the bidirectional transformer to generate entire samples. In our work, we propose using a bidirectional transformer diffusion, called Local Diffusion Transformer (LocDiT), to generate localized continuous token patches.
@@
Furthermore, the design also implies an inherent coarse-to-fine manner. Current approaches(Anastassiou et al., 2024; Du et al., 2024; Chen et al., 2024b) commonly explicitly delineate both coarse and fine features. A language model typically predicts the coarse feature, while another network achieves the transition from coarse to fine. However, multi-stage methods are prone to cumulative errors. In contrast, DiTAR functions in a seamless end-to-end manner, condensing each patch of tokens into an implicit lower-dimensional feature space that LocDiT subsequently expands into high-fidelity continuous tokens.
@@ 3.3 LM Guidance
Classifier-free guidance (CFG)(Ho & Salimans, 2022) is extensively employed to enhance the condition adherence of generative models. In diffusion, the unconditional and conditional models share parameters and are jointly trained by intermittently omitting the condition during training. At inference, the outputs from these two models are merged with a parameter $w$ balancing the trade-off between diversity and fidelity. This is equivalent to sampling under such a distribution $\tilde{p}_{\theta}\left(\bm{z}_{t},c\right)\propto p_{\theta}(\bm{z}_{t}|c)p_{\theta}(c|\bm{z}_{t})^{w}$, where $\theta$ denotes the model parameters, $c$ denotes the condition and $\bm{z}_{t}$ denotes the noisy sample at the time of $t$.
In the continuous-token LMs for conditional image generation, Li et al. (2024a) incorporates the class label into the diffusion head and applies CFG which significantly enhances performance. However, such an approach lacks generality. In DiTAR, we propose a more generalized approach: we place all conditions within the LM prefix inputs and only apply CFG to the diffusion decoder via the AR’s on-the-fly output. The $i$th output $\bm{h}_{i}$ of the AR essentially represents all historical inputs $(\bm{x}_{0},\bm{x}_{1},...,\bm{x}_{i})$. Specifically, we randomly replace the $\bm{h}_{i}$ with a null embedding along the sequence in training. At inference time, we samples from the distribution $p_{\theta}(\bm{z}_{i,t}|\bm{x}_{1},...,\bm{x}_{i-1})p_{\theta}(\bm{x}_{1},...,\bm{x}_{i-1}|\bm{z}_{i,t})^{w}$ by:
$\tilde{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{z}_{i,t},\boldsymbol{h}_{i})=(1+w)\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{i,t},\boldsymbol{h}_{i})-w\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{i,t})$ (2)
where $z_{i,t}$ denotes the $i$th noisy sample in the sequence at the time of $t$ for diffusion, $\epsilon_{\theta}$ denotes the score estimated by the LocDiT. Operating in the velocity space with a conditional flow-matching target is also equivalent. We show that the approach significantly enhances the model’s performance in Section 4.3.
@@ 3.4 Temperature for Continuous-Valued LMs
Temperature-based sampling, which strikes a balance between exploration and exploitation, is fundamental in discrete-valued LMs. However, its application in continuous-valued LMs has been less extensively studied. Li et al. (2024a) proposes a sampling method based on the DDPM reversed process(Ho et al., 2020), which is a first-order SDE solver, where the temperature is defined as the noise scaling factor at each step. Due to the stochastic nature of the Wiener process, SDE solvers require small step sizes, necessitating a higher number of steps for convergence(Song et al., 2020b). Additionally, this concept of temperature cannot be readily extended to the more widely used and faster ODE solvers prevalent in contemporary applications(Lu et al., 2022a, b; Song et al., 2020a). To address these issues, we propose a new sampling method with a new definition of temperature, which is compatible with the ODE solvers.
We define the temperature $\tau\in[0,1]$ as the time point to introduce noise while solving the reverse ODE of diffusion. Specifically, consider a Gaussian diffusion forward process defined on per-patch by $\bm{x}_{t}=\alpha_{t}\bm{x}_{0}+\sigma_{t}\bm{\varepsilon}$, where $\bm{x}_{0}\sim q_{data}(\bm{x}_{0})$, $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, $t\in[0,1]$, $\alpha_{t}$ and $\sigma_{t}$ collectively defines the flow path. At $\tau=1$, the sampling process is equivalent to the standard ODE sampling process, where we solve the reverse ODE $d{\bm{x}_{t}}=\bm{v}_{\theta}(\bm{x}_{t},t)dt$ with respect to $dt$ from $1$ to 00, and $\bm{x}_{1}\sim\mathcal{N}(\bm{0},\bm{I})$. At $\tau=0$, no random noise is introduced leading to a completely deterministic process. Considering that 00 is the value with the highest likelihood in standard Gaussian distribution, we define the greedy sampling by sampling from $x_{1}\equiv\bf{0}$, thus ensuring determinism.
When $10<\tau<$ 1, we introduce random noise at $\tau$ by using the forward process to diffuse an estimated $\bm{x}_{0}$. Eqn 3 and 4 summarize the iterative process, where the Euler method is used as the default ODE solver for illustration.
@@
where $\bm{x}_{\theta}$ represents the estimated $\bm{x}_{0}$, which can be derived from the estimated velocity or score through a linear transformation. The derivation process is detailed in Section A.3. In Section 4.3, we demonstrate that $\tau$ effectively balances diversity and stability. The sampling process is summarized in Algorithm 1.

### v3

@@ Abstract
Several recent studies have attempted to autoregressively generate continuous speech representations without discrete speech tokens by combining diffusion and autoregressive models, yet they often face challenges with excessive computational loads or suboptimal outcomes. In this work, we propose Diffusion Transformer Autoregressive Modeling (DiTAR), a patch-based autoregressive framework combining a language model with a diffusion transformer. This approach significantly enhances the efficacy of autoregressive models for continuous tokens and reduces computational demands. DiTAR utilizes a divide-and-conquer strategy for patch generation, where the language model processes aggregated patch embeddings and the diffusion transformer subsequently generates the next patch based on the output of the language model. For inference, we propose defining temperature as the time point of introducing noise during the reverse diffusion ODE to balance diversity and determinism. We also show in the extensive scaling analysis that DiTAR has superb scalability. In zero-shot speech generation, DiTAR achieves state-of-the-art performance in robustness, speaker similarity, and naturalness. [32]
[…]
@@ 1 Introduction
Recently, the autoregressive language model (LM) has demonstrated strong generative capabilities and scaling properties[1, 56, 60, 55], where discrete tokenization is commonly used. While discretization is natural for text, it is not as straightforward for other modalities, such as images, video, and audio. Due to bitrate limitations, discrete representation often fails to reconstruct complex modalities with high fidelity. In comparison, continuous tokens can better reconstruct the original information[15, 47]. To alleviate this challenge, a two-stage strategy is commonly applied for autoregressive multi-media generation, where a lossy discrete token is first generated by an LM, followed by a token-based diffusion for detail enrichment. However, such a cascaded design suffers from error accumulation and limits the scalability of the language model. Therefore, autoregressive modeling of continuous representations is meaningful.
@@
The diffusion model has been proven effective at modeling continuous representations[45, 47, 2], and integrating it with autoregressive models is a major research trend. [34] proposes incorporating a diffusion head into a language model for image generation, pioneering a new approach. However, the performance significantly falls short when a causal-attention language model forms the backbone. Another approach, such as ARDiT[38] or Transfusion[62], repurposes the language model’s parameters for diffusion, leading to substantial computational demands. This raises the question of whether we can autoregressively predict continuous tokens while ensuring high-quality generation and maintaining reasonable computational demands.
@@
To tackle this challenge, we introduce the Diffusion Transformer AutoRegressive (DiTAR) modeling, a framework that seamlessly combines transformer diffusion with a language model. We attribute the subpar performance of a language model with a diffusion head to the unidirectional dependency imposed by causal attention, which conflicts with the close inter-frame correlations characteristic of continuous tokens. In DiTAR, we propose a divide-and-conquer strategy that breaks continuous tokens into multiple patches. A language model is responsible for inter-patch prediction, while a diffusion transformer handles intra-patch prediction. Specifically, a diffusion transformer with bidirectional attention (DiT), noted for its state-of-the-art performance across various generative fields[37, 45, 2], is employed to predict localized patches, named LocDiT. Furthermore, we propose a broadly applicable guidance method and introduce historical contexts to LocDiT to further enhance its generative capabilities. After patchification, the causal language model handling long contexts receives shorter sequences, further reducing computational load.
@@
For inference, temperature-based sampling balances exploration and exploitation and is crucial in discrete-valued language models. However, Its application in continuous-valued LMs remains underexplored. We define temperature as the point at which noise is introduced along the reverse ODE trajectory and propose a temperature-based sampling method for the continuous-valued AR model. Distinct from previous SDE-based approaches[12, 34] that require numerous steps, our method suits quicker ODE solvers and adeptly balances diversity with determinism.
@@
DiTAR, rooted in a language model, excels at zero-shot generation tasks. We apply DiTAR to zero-shot text-to-speech (TTS) that aims to generate speech from unseen speakers’ prompts. Unlike systems[7, 8, 2, 13] using a coarse-to-fine pipeline, DiTAR simplifies the process by having the language model directly predict final features, obviating the need for multiple stages. It achieves state-of-the-art results in robustness, speaker similarity, and naturalness while demanding far less computational power than competing models.
@@ 2 Related Work
Integrating Autoregressive Language Model and Diffusion. Language models are primarily used for discrete representations, while diffusion excels in modeling continuous distributions. Integrating them for multimodal modeling is a crucial research direction. Some efforts [58, 38, 5] enable diffusion to have autoregressive capabilities by varying the denoising rates between consecutive tokens to achieve earlier predictions for preceding tokens. Transfusion[62] utilizes a shared transformer for both diffusion and language models, employing causal attention for discrete tokens and bidirectional attention for continuous tokens. However, it does not naturally support the autoregressive generation of continuous tokens. These approaches repurpose language model parameters for diffusion, which significantly increases computational demands as the sequence lengthens and the language model size grows. Most relevant to our work, Li et al. [34] proposes a diffusion head for next-token prediction. However, its application in causal language models results in relatively poor performance.
@@
Patchification in Generative Modeling. In speech[57, 43, 7], image[45, 35], and video generation[37], the patchification technique is widely applied. In these works, patchification primarily aims to reduce the computational load by shortening the sequence length. However, in this paper, patchification not only lowers computational demands but also enables bidirectional modeling on patches within our autoregressive framework, further improving modeling effectiveness.
@@
Zero-Shot Text-to-Speech. Zero-shot TTS aims to generate speech with prompts from unseen speakers. Existing works can be divided into two categories: multi-stage and single-stage. Multi-stage approaches[7, 8, 31, 23, 22, 2] represent speech using various representations, primarily divided into coarse and fine categories. The autoregressive language model is often adopted to predict the coarse representations, usually in discrete values and low information, such as semantics[26, 2, 21, 10, 4] and prosody[23, 22, 24]. And then another model is adopted to conduct coarse-to-fine. Single-stage methods focus on generating high-information continuous representations, such as Mel spectrograms or latents of auto-encoders, which can directly reconstruct audio waveforms. Diffusion based on the non-causal attention transformer is the primary method employed[9, 14, 16, 29]. Additionally, some approaches[43, 57, 50, 33] directly use AR to model Mel-spectrograms of speech but often involve using Dropout[53] at the input stage to guarantee generation robustness, resulting in weaker in-context learning capabilities. These methods are better suited for in-set speaker TTS. This paper introduces a single-stage autoregressive method for zero-shot TTS, achieving SOTA generation robustness and voice similarity.
3 Approach
3.1 Overview
@@ 3.1.1 Formulation
DiTAR is an autoregressive model via next-token prediction. Consider a sequence of continuous tokens $\bm{x}=(\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{N})$. We can factorize the joint distribution of the sequence by the chain rule of probability:
@@
We validate the effectiveness of DiTAR on the zero-shot text-to-speech task. Based on the formulation, we regard zero-shot TTS as a conditional continuation task for the AR model like[7], where prompting texts, target text, and prompting speech are concatenated and fed into the model as prefix context, then the model autoregressively generates the target speech given the context.
@@ 3.1.2 Overall architecture
Li et al. [34] finds that the performance of the causal-attention autoregressive model combined with diffusion loss is significantly inferior to systems utilizing full attention. To address the issue, we propose a divide-and-conquer strategy, where a long sequence of continuous tokens is divided into multiple patches. A language model is responsible for inter-patch prediction, while a diffusion transformer handles intra-patch prediction. As shown in Figure 1, the backbone of our system is a causal-attention transformer with next-token prediction. Each patch of continuous tokens is processed with an aggregation encoder into a single vector, which is then fed into the AR model to get the output embedding $\bm{h}_{t}$. $\bm{h}_{t}$ serves as the condition of the following diffusion decoder, LocDiT. Following [34], a diffusion loss is used for the output continuous tokens at training time.
@@ 3.2 LocDiT: Next-Patch Bidirectional Modeling
The diffusion transformer[45, 37] has achieved success across numerous generative fields. These approaches leverage the full receptive field of the bidirectional transformer to generate entire samples. In our work, we propose using a bidirectional transformer diffusion, called Local Diffusion Transformer (LocDiT), to generate localized continuous token patches.
@@
Furthermore, the design also implies an inherent coarse-to-fine manner. Current approaches [2, 13, 7] commonly explicitly delineate both coarse and fine features: a language model typically predicts the coarse feature, while another network achieves the transition from coarse to fine. However, multi-stage methods are prone to cumulative errors. In contrast, DiTAR functions in a seamless end-to-end manner, condensing each patch of tokens into an implicit coarse feature space that LocDiT subsequently expands into high-fidelity continuous tokens.
@@ 3.3 LM Guidance
Classifier-free guidance (CFG)[19] is extensively employed to enhance the condition adherence of generative models. In diffusion, the unconditional and conditional models share parameters and are jointly trained by intermittently omitting the condition during training. At inference, the outputs from these two models are merged with a parameter $w$ balancing the trade-off between diversity and fidelity. This is equivalent to sampling under such a distribution $\tilde{p}_{\theta}\left(\bm{z}_{t},c\right)\propto p_{\theta}(\bm{z}_{t}|c)p_{\theta}(c|\bm{z}_{t})^{w}$, where $\theta$ denotes the model parameters, $c$ denotes the condition and $\bm{z}_{t}$ denotes the noisy sample at the time of $t$. For discrete-valued language models, classifier-free guidance[49] often involves computing the language model twice to obtain conditional and unconditional logits, which can be computationally expensive.
For DiTAR, a model that incorporates both an LM and a diffusion head, we propose LM guidance, an effective approach that requires only two computations of the diffusion head and one computation of the LM. Specically, the $i$th output $\bm{h}_{i}$ of the LM essentially represents all historical inputs $(\bm{x}_{0},\bm{x}_{1},...,\bm{x}_{i})$ and we randomly replace the $\bm{h}_{i}$ with a dummy embedding $\boldsymbol{h}_{\varnothing}$ along the sequence in training. At inference time, we samples from the distribution $p_{\theta}(\bm{z}_{i,t}|\bm{x}_{1},...,\bm{x}_{i-1})p_{\theta}(\bm{x}_{1},...,\bm{x}_{i-1}|\bm{z}_{i,t})^{w}$ by:
$\tilde{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{z}_{i,t},\boldsymbol{h}_{i})=(1+w)\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{i,t},\boldsymbol{h}_{i})-w\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{i,t},\boldsymbol{h}_{\varnothing})$ (2)
where $\boldsymbol{z}_{i,t}$ denotes the $i$th noisy sample in the sequence at the time of $t$ for diffusion, $\boldsymbol{\epsilon}_{\theta}$ denotes the score estimated by the LocDiT, $\boldsymbol{h}_{\varnothing}$ denotes the dummy vector representing unconditional modeling. Operating in the velocity space with a conditional flow-matching target is also equivalent. We show that the approach significantly enhances the model’s performance in section 4.4.
@@ 3.4 Temperature for Continuous-Valued LMs
Temperature-based sampling, which strikes a balance between exploration and exploitation, is fundamental in discrete-valued LMs. However, its application in continuous-valued LMs has been less extensively studied. Li et al. [34] proposes a sampling method based on the DDPM reversed process[20], which is a first-order SDE solver, where the temperature is defined as the noise scaling factor at each step. Due to the stochastic nature of the Wiener process, SDE solvers require small step sizes, necessitating a higher number of steps for convergence[52]. Additionally, this concept of temperature cannot be readily extended to the more widely used and faster ODE solvers prevalent in contemporary applications[40, 41, 51]. To address these issues, we propose a new sampling method with a new definition of temperature, which is compatible with the ODE solvers.
We define the temperature $\tau\in[0,1]$ as the time point to introduce noise while solving the reverse ODE of diffusion. Specifically, consider a Gaussian diffusion forward process defined on per-patch by $\bm{x}_{t}=\alpha_{t}\bm{x}_{0}+\sigma_{t}\bm{\varepsilon}$, where $\bm{x}_{0}\sim q_{data}(\bm{x}_{0})$, $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, $t\in[0,1]$, $\alpha_{t}$ and $\sigma_{t}$ collectively defines the flow path. At $\tau=1$, the sampling process is equivalent to the standard ODE sampling process, where we solve the reverse ODE $d{\bm{x}_{t}}=\bm{v}_{\theta}(\bm{x}_{t},t)dt$ from $1$ to 00, where $\bm{x}_{1}\sim\mathcal{N}(\bm{0},\bm{I})$ and $\bm{v}_{\theta}$ denotes predicted vector field. At $\tau=0$, no random noise is introduced, leading to a completely deterministic process. Considering that 00 is the value with the highest likelihood in the standard Gaussian distribution, we define the greedy sampling by sampling from $x_{1}\equiv\bf{0}$, thus ensuring determinism.
When $10<\tau<$ 1, we introduce random noise at $\tau$ by using the forward process to diffuse an estimated $\bm{x}_{0}$. Eq 3 and 4 summarize the iterative process, where the Euler method is used as the default ODE solver for illustration.
@@
where $\bm{x}_{\theta}$ represents the estimated $\bm{x}_{0}$, which can be derived from the estimated velocity or score through a linear transformation. The derivation process is detailed in section 6.3. In section 4.4, we demonstrate that $\tau$ effectively balances diversity and stability. The sampling process is summarized in Algorithm 1.
Algorithm 1 Temperature sampling
  $\eta\leftarrow\underset{n=1,2,...,N}{\operatorname{argmin}}|t_{n}-\tau|$
  if $\eta=N$ then
     Sample initial noise $\bm{x}_{t_{N}}\sim\mathcal{N}(\bm{0},\bm{I})$
  else
     initial noise $\bm{x}_{t_{N}}\leftarrow\bm{0}$
  end if
  for $n=N-1$ to $1$ do
     $\hat{\bm{v}}\leftarrow\bm{f_{\theta}}(\bm{x}_{t_{N}},t_{N})$
     if $t_{n}=\eta$ then
        $\hat{\bm{x}}\leftarrow\mathcal{F}(\hat{\bm{v}})$
        Sample $\bm{z}\sim\mathcal{N}(\bm{0},\bm{I})$
        $\bm{x}\leftarrow\alpha_{t_{n+1}}\hat{\bm{x}}+\beta_{t_{n+1}}\bm{z}$
     else
        $\bm{x}\leftarrow\bm{\Psi}(\hat{\bm{v}},n+1,n)$
     end if
  end for
  Output: $\bm{x}$
3.5 Implemenations
3.5.1 Continuous speech tokenization
Following LDM[47], we use a variational auto-encoder (VAE)[27] to convert the waveform into the distribution of latent $z$, represented by mean and variance. The encoder of the VAE consists of multiple layers of the convolutional network, and the decoder’s architecture follows BigVGAN[30]. The adversarial training scheme also follows Lee et al. [30]. We adopt the multi-period discriminator (MPD) and multi-scale discriminator (MSD) proposed by Kong et al. [28] as our discriminators. In our setup, the 24000hz waveform is compressed into 40Hz latent with a dimension of 64.
3.5.2 Model
DiTAR consists of three modules: aggregation encoder, language model, and decoder (LocDiT). In our implementation, all of them are based on a transformer architecture. Specifically, both the encoder and decoder employ bidirectional attention masks, while the LM utilizes a causal attention mask. All transformers adopt the Pre-Norm[59] architecture and utilize RMSNorm[61] and RoPE [54]. Each patch of continuous tokens, together with a learnable special token positioned at the beginning of the sequence similar to [11], is fed into the aggregation encoder. The output corresponding to the special token’s position serves as the aggregation embedding. Aggregation embeddings from different patches form a sequence that is processed by the LM. In LocDiT, LM’s outputs and the time embedding are added, along with historical context patches and noisy target tokens, forming a new sequence that serves as the input of LocDiT. When calculating the loss, we only consider the output corresponding to the position of noisy target tokens. During training, LM’s output is randomly replaced by a vector of all zeros with a probability of 0.1 to enable LM guidance for LocDiT.
3.5.3 Diffusion formulation
Following Song et al. [52], Lu and Song [39], we adopt a variance-preserving diffusion process defined by
$\bm{x}_{t}$ $=\alpha_{t}\bm{x}_{0}+\sigma_{t}\bm{\varepsilon}$ (5)
where $\bm{x}_{0}\thicksim q(\bm{x}_{0})$ denotes the data, $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ denotes the standard Gaussian noise, $t\in[0,1]$. We employ a conditional flow-matching loss[36]:
where the velocity is defined as: $\bm{v}(\bm{x}_{t},t)=\dot{\bm{x}_{t}}=\dot{\alpha}_{t}\bm{x}_{t}+\dot{\sigma}_{t}\bm{\varepsilon}$. At inference time, We employed the DDIM sampler[51], which is essentially an Euler ODE sampler with respect to signal-to-noise ratio $\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}}$ instead of $t$, proved to better align with the semi-linear property of the diffusion ODE[40].
3.5.4 Zero-shot TTS system
The text sequence is converted into phonemes and processed through a lookup table to obtain text embeddings. Speech tokens are processed by the aggregation encoder to produce speech embeddings, which are then concatenated with the text embeddings. The embedding sequence serves as the input of the LM of DiTAR. During training, the loss of the text is not included. Additionally, we introduce a binary classifier with a fully connected layer at LM’s output to predict when to stop following Wang et al. [57], Li et al. [33]. The loss function for zero-shot TTS can be summarized as $L=L_{diff}+L_{\text{stop}}$. During inference, text, target text, and prompting audio are fed as prefix input to DiTAR’s LM, which then autoregressively generates the target audio.
4 Experiments
4.1 SOTA Performance in Zero-Shot TTS
4.1.1 Setup
To ensure a fair evaluation of zero-shot TTS, it is essential to consider prompt audio, texts, and tools. We standardize these variables to facilitate a more objective and fair comparison between systems.
Training and Evaluation Dataset. We consider two open-source datasets as our training dataset. 1) Librilight[25], containing 60K hours of English speech data from LibriVox audiobooks. 2) Emilia[18], a multilingual dataset containing around 100k hours of speech data.
We adopt three open-source datasets for evaluation: 1) LibriSpeech(PC)[44, 42] test-clean, containing 40 distinct English speakers and a 5.4-hour speech. We employ two established subsets: subset A from NaturalSpeech3, featuring 40 three-second speech prompts and 40 target samples, and subset B from F5TTS, which includes 40 prompts and 1127 samples. 2) Seed-ZH: a subset from DiDiSpeech 2[17], a Chinese speech dataset, containing 1088 prompts and targets. 3) Seed-EN: a subset from Common Voice[3], a crowdsourcing English speech dataset with diverse accents, containing 2020 prompts and targets.
Evaluation Metrics. We evaluate four objective metrics: 1) Word Error Rate (WER), which assesses generation robustness. For consistency, we use the same ASR setups as previous studies to transcribe generated speech across different test sets. Specifically, we adopt a Hubert-based model [^1] https://huggingface.co/facebook/hubert-large-ls960-ft and Faster-whisper-large-v3  [^2] https://huggingface.co/Systran/faster-whisper-large-v3,version:0.10.1[46] for the subset A and B of Librispeech test-clean, respectively, Whisper-large-v3333https://huggingface.co/openai/whisper-large-v3 for Seed-EN and Paraformer-zh for Seed-ZH. 2) Speaker similarity to the prompt audio (SIM). Specifically, we employed WavLM-large[6] to compute the cosine distance between the generated and reference speech. 3) UTMOS[48], an automatic predicted mean opinion score (MOS) to evaluate speech quality. 4) Floating point operations (FLOPs), measuring the computational load of models. The calculation process is detailed in 6.4.
For subjective evaluation, we employ four MOS metrics: 1) N-MOS for naturalness, 2) Q-MOS for sound quality, 3) S-MOS for speaker voice similarity 4) CMOS for side-by-side comparison with human audios.
Model Setup and Baselines. We benchmark DiTAR against diverse zero-shot TTS systems with varying architectures, including both multi-stage and single-stage schemes that operate in discrete or continuous value spaces. We conduct comparisons using DiTAR with 0.6 billion parameters and a patch size of 4. During inference, DiTAR’s LocDiT uses an NFE (Number of Function Evaluations) of 10. Specific details about the parameters of DiTAR are provided in Appendix 6.2.
4.2 Scaling Behaviors
The language model and diffusion decoder benefit more from scaling. We scale the encoder, language model, and decoder (LocDiT) individually to assess their impacts, starting with a 0.4B parameter model trained on 280k hours data . Given that the encoder and decoder process shorter sequences, we allocated more parameters to the language model empirically. Table 4 demonstrates that enlarging the language model and LocDiT enhances performance, whereas increasing the encoder size has little effect on outcomes.